Assignment Question(s):

Q1-

Abdulaziz Co. purchased a machine in 2013 for 50,000 that has a useful life of 5 years with a salvage value of 5,000.

Calculate the depreciation expense, accumulated depreciation, book value throughout its useful life using:

1- Straight-line Method.

2- Units of Production Method if the machine produces 100,000 units.

Here is a table of units produced each year:

First

Second

Third

Fourth

Fifth

23,000

25,000

-

30,000

22,000

3- Double Declining Balance Method .

Q2- On June 1, 2019, ABC Company signed a $25,000, 120-day, 6% note payable to cover a past due account payable.
a. What is the total amount of interest to be paid on this note?
b. Prepare ABC Company's general journal entry to record the issuance of the note payable
c. Prepare ABC Company's general journal entry to record the payment of the note on
September 29, 2019.

Q3. What are the characteristics of corporations ?
Answer #1

Q1.
1 - Straight line method
Depreciation = (Original Value - Salvage Value) / Useful Life
= ($50000 - 5000) / 5 = $9000 per year

Year Beginning Book Value Depreciation Accumulated Depreciation Ending Book Value
1 $                50,000 $             9,000 $                 9,000 $               41,000
2 $                41,000 $             9,000 $               18,000 $               32,000
3 $                32,000 $             9,000 $               27,000 $               23,000
4 $                23,000 $             9,000 $               36,000 $               14,000
5 $                14,000 $             9,000 $               45,000 $                 5,000

2. Units of Production Method
Depreciation = (Original Value - Salvage Value) / Total Production
= ($50000-5000) / 100000 = $0.45 per unit

Year Beginning Book Value Depreciation Accumulated Depreciation Ending Book Value
1 $                50,000 $           10,350 $               10,350 $               39,650
2 $                39,650 $           11,250 $               21,600 $               28,400
3 $                28,400 $                    -   $               21,600 $               28,400
4 $                28,400 $           13,500 $               35,100 $               14,900
5 $                14,900 $             9,900 $               45,000 $                 5,000

3. Double Declining Balance method
Depreciation = Beginning Value x 2 times straight line rate
Straight line rate = 1/5x 100 = 20%

Year Beginning Book Value Depreciation Accumulated Depreciation Ending Book Value
1 $                50,000 $           20,000 $               20,000 $               30,000
2 $                30,000 $           12,000 $               32,000 $               18,000
3 $                18,000 $             7,200 $               39,200 $               10,800
4 $                10,800 $             4,320 $               43,520 $                 6,480
5 $                   6,480 $             1,480 $               45,000 $                 5,000


Depreciation for Year 5 is lower than 40% of Beginning book value, as ending value cannot be lower than salvage value

As per HOMEWORKLIB RULES we are supposed to answer 1 question completely, so please post other questions separately